OpenStudy (anonymous):

PLEASE HELLLLLP! Which of the following ordered pairs is not a solution to the inequality y greater than or equal to -4x - 2? (-1, 0) (0, 0) (4, 4) (1, 3)

OpenStudy (anonymous):

Plug the numbers in for the equation \[y \ge-4x-2 \] The coordinates in the answers are in the form\[(x,y)\] Plug and solve

OpenStudy (anonymous):

I am so confused

OpenStudy (anonymous):

@Mike546

OpenStudy (anonymous):

Ok the first number of the ordered pair is x and the second number is y. Plug those values into the inequality and see if any of them make the inequality true.

OpenStudy (anonymous):

ok so is it (4, 4)?

OpenStudy (anonymous):

Why do you think that? Please explain your reasoning.

OpenStudy (anonymous):

is it wrong?

OpenStudy (anonymous):

I am not saying that I just want to see your reasoning...or are you just guessing?

OpenStudy (anonymous):

no I pluged it in and I got (o,0)

OpenStudy (anonymous):

I am not sure what you mean. Let's look at (4,4)...If you put 4 in for x into the right side of the inequlity... -4x-2 what is the result?

jhonyy9 (jhonyy9):

y>= -4x-2 (-1, 0) x= -1 so this mean y>= -4(-1)-2 =5-2 =3 so but there added that y=0 from this result that not is right (0, 0) (4, 4) (1, 3) do you like continue it for meadl ?

jhonyy9 (jhonyy9):

medal -- sorry

jhonyy9 (jhonyy9):

@cambrige camon ,thisis easy ,courage !!!